OpenStudy (anonymous):

Simplify A. 4 B. 7.2 C. 8 D. 36

OpenStudy (anonymous):



OpenStudy (anonymous):

4

OpenStudy (anonymous):

before the carrot?what does that mean

OpenStudy (anonymous):

thats not one of the answers thats up on the list

OpenStudy (anonymous):

258

OpenStudy (anonymous):

Not very sure, sorry for wasting you're time :/

OpenStudy (anonymous):

no its ok.

OpenStudy (chrisdbest):

It's C: 8

OpenStudy (anonymous):

can you show me how?

OpenStudy (anonymous):

Same :P

OpenStudy (freckles):

\[\sqrt[4]{16}=2 \text{ since } 2^4=(2)(2)(2)(2)=16\]